OpenStudy (anonymous):

What is the absolute value equation l 5x - 10 l= 20?

jigglypuff314 (jigglypuff314):

l 5x - 10 l= 20 can be separated into 5x - 10 = 20 and -(5x - 10) = 20 then solve for x -5x + 10 = 20

OpenStudy (anonymous):

oh okay thank you